OpenStudy (anonymous):

3^{2x}-3^{x+1}+8/9=0

OpenStudy (anonymous):

\[3^{x}=t\]

OpenStudy (anonymous):

\[3^{2x}-3^{x}.3^{1}+8/9=0\]

OpenStudy (anonymous):

\[t ^{2}-3t+8/9=0\]

OpenStudy (anonymous):

hope you can solve now.....

OpenStudy (anonymous):

use nitz method!1

OpenStudy (anonymous):

0.45 -3.45 i got it.

OpenStudy (anonymous):

is it rite?? @nitz

OpenStudy (anonymous):

that is t now subs t = 3^x

OpenStudy (anonymous):

i am getting t=1/3 and t=8/3.

OpenStudy (anonymous):

how can i slove 8/9 -3+-\[-3+-\sqrt{3^{2}}+4*1*8/9/2*1\]

OpenStudy (anonymous):

what is it?????

OpenStudy (anonymous):

\[-3+-\sqrt{9+32/9}/2\]

OpenStudy (anonymous):

\[9t ^{2}-27t+8=0\]

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

this is the required quadratic i am getting........ it can be factorised into linear factors

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I GOT 1/3 and 8/3

OpenStudy (phi):

the quadratic formula is \[ t= \frac{1}{2}(-b ± \sqrt{b^2-4ac} ) \] for your problem \[ t^2- 3 t +\frac{8}{9} =0 \] you get \[ t= \frac{1}{2}(-(-3) ± \sqrt{9-\frac{32}{9}}) \]

OpenStudy (anonymous):

@nitz thanks

OpenStudy (anonymous):

i m confused which answer is right @nitz and @phi

OpenStudy (phi):

that simplfies to \[ t= \frac{1}{2}(3±\sqrt{\frac{49}{9}} ) \] or \[ t= \frac{3}{2} ± \frac{7}{6} = \frac{16}{6} or \frac{2}{6}\]

OpenStudy (phi):

which simplifies to 8/3 and 1/3

OpenStudy (phi):

x= 3^(8/3) or x= 3^(1/3)

OpenStudy (anonymous):

why is t=1/2

OpenStudy (phi):

I should have typed \[ t= \frac{1}{2a}(-b ± \sqrt{b^2-4ac} ) \] which is the quadratic formula

OpenStudy (anonymous):

@phi why dont you directly factorise.... \[9t ^{2}-27t+8=0\] \[9t ^{2}-24t-3t+8=0\] \[3t(3t-8)-1(3t-8)=0\] (3t-1)(3t-8)=0 t=1/3 and t=8/3 then...\[3^{x}=1/3 and 3^x=8/3

OpenStudy (anonymous):

is not a 1 quadratic formula is \[-b+-\sqrt{b2-4*a*c}/2*a\]

OpenStudy (phi):

yes, but remember the -b is also divided by 2a

OpenStudy (phi):

but t is not 1/2 . t is 1/3 or 8/3 as for how to solve this, use whatever method you are most comfortable with. all variations: quadratic formula or factoring give the same answer.

OpenStudy (anonymous):

\[9t ^{2}-27t+8=0\] Use factorization: \[9t^2 - 3t - 24t + 8 = 0\] \[3t(3t-1) - 8(3t - 1) = 0 \implies (3t-1)(3t-8) = 0\] \[t = \frac{1}{3} \qquad Or \qquad t = \frac{8}{3}\]

OpenStudy (anonymous):

\[3^x = \frac{1}{3} \implies 3^x = 3^{-1} \implies x = -1\]

OpenStudy (anonymous):

\[3^x = \frac{8}{3} \implies 3^{x+1} = 8\] taking log(10): \[\large (x+1) \cdot \log_{10}(3) = 3 \cdot \log_{10}(2)\] \(log_{10} (3) = 0.4771 \) and \(log_{10}(2) = 0.3010\) Put this values and solve for x now..

OpenStudy (anonymous):

yupp..thanks

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

:)